Aggregator

Operator SIGN IN

2 days 3 hours ago
Log Number: 4467928Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Anthony
DATE:             01/30/2026
SHIFT:               Swing

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log
anthony

COMMENT ON RF work for 2026 HCO

2 days 3 hours ago
Log Number: 4467927Logbook(s): Body: 

Several zones in the North and South linacs have Sorenson power supplies that provide current to the C100 solenoids. These power supplies need maintenance every year due to getting hot and shutting down. We currently have six and possibly 7 in the North Linac zones 1L24 and 1L25 that need pulled and cooling fans checked/replaced. These two zones wouldn't go into HV for PSS certification this week and will not be ready until the supplies are repaired and reinstalled so that we can go into HV in the zone. Larry Farrish has 6 in for repair that hopefully will work early next week. Kerns 7283.

kerns

COMMENT ON Investigate Trim Rack Problems Found doing HCO

2 days 3 hours ago
Log Number: 4467925Logbook(s): Body: 

We had to re-connect a comm cable inside the Utility Chassis for E301B11 Trim Rack, all magnets where set to +/- low current for 30 min and then cycled. All magnets and Trim Rack are ready for operation.

We had to replace a neg. bootstrap fuse on the back of the Utility Chassis in W301B13 Trim Rack. all magnets where set to +/- low current for 30 min and then cycled. All magnets and Trim Rack are ready for operation.

We had to replace a front door switch in E501B11 and a back door switch in E501B12 Trim Rack. all magnets where set to +/- low current for 30 min and then cycled.

diggs

COMMENT ON Wien filter chiller relocation and H-Wien cooling plates installation

2 days 4 hours ago
Log Number: 4467924Logbook(s): Body: 

Thermocouple wire terminations have been completed today by Kevin with commercial ends wrapped in Kapton. Kevin installed a connector block where the TC wires coming from the Labjack T7 transition to the TC terminations. The V-Wien TCs are taped on top (1I04 Ch. 1) and bottom (1I04 Ch. 2) of the magnet yoke. The H-Wien TCs are taped on the magnet yoke beam right (1I06 Ch. 1) and beam left (1I06 ch. 2).

PVs for these channels are now in EPICS reporting in Celsius, and are being archived: https://logbooks.jlab.org/entry/4467142

chgarcia

HYST AREA 7-0 Cycles Magnets

2 days 4 hours ago
Log Number: 4467923Logbook(s): Body: 

HYST AREA Version 7-0 has successfully cycled the following:

MQA7S04
MBC9S02H
MBC9S02V
MQA3S04
MBDBS01H
MBDBS01V
MBT1S05H
MBT1S05V
MQB1S06
MQD9S02
MQA9S02
MQC9S00
MBC7S03V
MQA9S01
MBC9S01H
MBC9S01V

Total Time: 1.933 mins

diggs

COMMENT ON Vacuum Activities CEBAF 2/18/2025 through 2/18/26

2 days 4 hours ago
Log Number: 4467922Logbook(s): Body: 

2L24. Opened downstream cold valve. Closed up and reestablished insulating vacuum in preparation for leak check. Cleanroom, turbo pump and equipment removed from upstream of 25.

2L25. Upstream cold valve open. New vat valves installed at VBV2L25B & VBV2L26A. Final leak check was good. Minimal detectable leak 2e-10. 2L25 downstream & 2L26 upstream cold valves will remain closed through the weekend. Currently have VIP2L26A & VIP2L26B on local pig power supplies.
Monitored/pumping 2L26 cavity vacuum from VIP2L27A.

fhumphry

COMMENT ON Upgrade 2L03 to a C75 zone.

2 days 4 hours ago
Log Number: 4467920Logbook(s): Body: 

Initial LLRF test in to shorts is followed for 2L03. HPA interlocks were checked and the Klystron characterizations are followed.Max Kly powers :
1)4.835 kW
2)3.556 kW
3)4.737 kW
4)4.276 kW
5)5.100 kW
6)4.656 kW
7)Bypassed
8)3.821 kW

-plawski, bachiman, jkt

jkt

HYST AREA 7-0 Cycles Magnets

2 days 4 hours ago
Log Number: 4467918Logbook(s): Body: 

HYST AREA Version 7-0 has successfully cycled the following:

MQB2T03
MBC2T03H
MBC2T03V
MQA4T03
MBC4T03H
MBC4T03V
MQA6T02
MBCAT02H
MBCAT02V
MBC6T01V
MBCAT01V
MBC8T01H
MBC8T01V
MBC2T02H
MQA4T02
MBCAT01H
MBC2T01H
MBC2T01V
MQA4T01
MBC4T01H
MBC4T01V
MQA6T01
MBC6T01H

Total Time: 2.500 mins

banks

BSList #117047: more data

2 days 4 hours ago
Log Number: 4467914Logbook(s): Entry Makers: brukerBody: 

Simulations predict one should be able to find the smallest beam size in the VWien by scanning MFX2I01/1I03 and setting MFX0I01 to three different values for each of these settings and for Wien on/off. Why three values: It doesn't need a full lens scan as you would do it for phase-space reconstruction, but a single MFX0I01 setting isn't enough because MFG1I04A/B is not a thin lens; the beam size at the harp is ambiguous. It would be unambiguous with MFG1I04A/B off, but this is hopeless because of the DP aperture.

bruker

HYST AREA 7-0 Cycles Magnets

2 days 5 hours ago
Log Number: 4467913Logbook(s): Body: 
HYST AREA Version 7-0 has FAILED to cycle:

MQC2A05 -- W202B17 -- wa2mag -- Failed to set after 2 cycle attempts
MQA4A04 -- W202B17 -- wa2mag -- Failed to set after 2 cycle attempts
MQA6A04 -- W202B17 -- wa2mag -- Failed to set after 2 cycle attempts
MBC6A04V -- W202B17 -- wa2mag -- Failed to set after 2 cycle attempts
MBT2A05V -- W202B17 -- wa2mag -- Failed to set after 2 cycle attempts
MQA8A04 -- W202B17 -- wa2mag -- Failed to set after 2 cycle attempts
MBC4A04V -- W202B17 -- wa2mag -- Failed to set after 2 cycle attempts
banks

APPROVED YBBS06 Leak Repair

2 days 5 hours ago
Log Number: 4467910Logbook(s): Body: 

We're in the process of coordinating the timing of this work. Will update ATLis with scheduling details as they solidify. Be sure the PTW is on-issue before proceeding.

moser

APPROVED RF work for 2026 HCO

2 days 5 hours ago
Log Number: 4467908Logbook(s): Body: 

Looks good. For each job before proceeding please be sure that...
1) the relevant PTW is on-issue
2) the work has been coordinated with the MCC
3) there is a comment on this ATLis regarding what is about to be done

Thanks.

moser

Adjusted Wien chiller temperature set point, and a few observations on temperature

2 days 5 hours ago
Log Number: 4467903Logbook(s): Body: 

With the chiller temperature set to 18 Celsius, the V-Wien negative HVPS tripped off within 56 hours after being set for +90 deg spin rotation.
Preliminary observations show that the elapsed time between trip offs seems to be longer as the chiller temperature is decreased. The pressure in the Wien vacuum chamber increases with temperature, leading to HVPS trips that result in beam line pressure spikes into the 100 micro-A range (was reported by the ion pump controllers).

chgarcia

Autolog save

2 days 5 hours ago
Log Number: 4467892Logbook(s): Tags: Body: saver ran from opsl06

Saved: bruker 01-30-26 13:31:44 {A:5 B:3 C:1 E:340 A=Tune B=BS C=BS D=BS 0R08=0.00 0R08 dump in. A dump(s) in.
bruker

Follow-up Re: Summary of Laser Table Work

2 days 6 hours ago
Log Number: 4467884Logbook(s): Entry Makers: suleimanBody: 

Controlled access to CEBAF laser enclosure to investigate bad Moller PITA scans.

-First problem found was that RTP Box 2 was being erroneously triggered with nHelicity. Riad fixed that.

-Second problem found was misconnection of RTP Box 2's outputs

-Per the Crystal Graphic, Box 2's positive output (labeled V3+) is supposed to be connected to the bottom of crystal 2, and Box 2's negative output (labeled V4-) is supposed to be connected to the top of crystal 2. (See Figure 1)

kortze

HYST AREA 7-0 Cycles Magnets

2 days 6 hours ago
Log Number: 4467883Logbook(s): Body: 
HYST AREA Version 7-0 has FAILED to cycle:

MQC2A05 -- W202B17 -- wa2mag -- Failed to set after 2 cycle attempts
MQA4A04 -- W202B17 -- wa2mag -- Failed to set after 2 cycle attempts
MQA6A04 -- W202B17 -- wa2mag -- Failed to set after 2 cycle attempts
MBC6A04V -- W202B17 -- wa2mag -- Failed to set after 2 cycle attempts
MBT2A05V -- W202B17 -- wa2mag -- Failed to set after 2 cycle attempts
MQA8A04 -- W202B17 -- wa2mag -- Failed to set after 2 cycle attempts
MBC4A04V -- W202B17 -- wa2mag -- Failed to set after 2 cycle attempts
diggs

Shift Information

Shift Day
CEBAF Program
MCC Crew Chief Humphry
MCC Operators
Program Deputy Mock
Announcements

Access Status

Recent PSS History

recent pss history graph